# A Cognitive Paradigm Approach to Probe the Perception-Reasoning Interface in VLMs

### v3

@@ Abstract
A fundamental challenge in artificial intelligence involves understanding the cognitive processes underlying visual reasoning in sophisticated models like Vision-Language Models (VLMs). How do these models integrate visual perception with abstract thought, especially when reasoning across multiple images? Drawing inspiration from cognitive science, this paper introduces a structured evaluation framework using Bongard Problems (BPs) – a classic test of visual abstraction – to dissect the perception-reasoning interface in VLMs. We propose three distinct evaluation paradigms, mirroring human problem-solving strategies: Direct Visual Rule Learning (DVRL; holistic processing), Deductive Rule Learning (DRL; rule extraction and application), and Componential Analysis (CA; analytical decomposition via textual descriptions). These paradigms allow us to systematically vary the cognitive load and probe specific processing stages. Notably, the CA paradigm enables the evaluation of multi-image reasoning even in VLMs architecturally limited to single images and facilitates the isolation of reasoning capabilities from perceptual limitations by controlling the descriptive input. Ablation studies further confirm that reasoning abilities improve significantly when perceptual challenges are mitigated. Our framework provides a valuable diagnostic tool, highlighting the need to enhance visual processing fidelity for achieving more robust and human-like visual intelligence in AI.
A Cognitive Paradigm Approach to Probe the Perception-Reasoning Interface in VLMs
Mohit Vaishnav1,2and Tanel Tammet1 1Applied Artificial Intelligence Group, Tallinn University of Technology, Estonia 2Kimova AI, Tallinn, Estonia mohit.vaishnav@taltech.ee, tanel.tammet@taltech.ee
@@ 1 Introduction
Human cognition seamlessly integrates rich visual perception with abstract reasoning, enabling us to understand complex scenes, infer relationships, and solve problems grounded in the visual world Kunda (2020); Lake et al. (2017). Replicating this visual intelligence in artificial systems remains a central goal of AI research Cao et al. (2024). While Vision-Language Models (VLMs) represent significant progress Zhao et al. (2023); Radford et al. (2021), the internal mechanisms by which they perform visual reasoning – particularly tasks requiring abstraction and comparison across multiple inputs – are still poorly understood. How do these models bridge the gap between pixel-level input and conceptual understanding, and how do their strategies compare to human cognitive processes?
To explore these questions, we turn to Bongard Problems (BPs) Bongard (1968). BPs present a unique challenge: identifying an abstract visual rule that distinguishes a set of positive image examples from a set of negative ones. Solving BPs demands not only accurate perception of individual images but also the ability to identify shared properties, ignore irrelevant features, formulate an abstract rule, and apply it to novel instances – processes central to human concept learning and inductive reasoning Hofstadter (1995). Furthermore, BPs inherently require processing and comparing multiple images, a capability many standard VLM architectures handle implicitly or not at all Wu et al. (2024); Jiang et al. (2022).
This paper introduces an evaluation framework explicitly designed to probe the cognitive processes underlying VLM performance on such challenging visual reasoning tasks, using natural image BPs Wu et al. (2024) as our primary testbed. Our core contribution is an evaluation methodology grounded in cognitive science principles. We propose three distinct paradigms that structure the BP task differently, mirroring established human problem-solving strategies:
@@
    Direct Visual Rule Learning (DVRL): Simulates holistic or gist-based processing Biederman (1987), requiring the model to analyze all images simultaneously.
@@
    Deductive Rule Learning (DRL): Mimics explicit, rule-based deduction Rips (1994), separating rule extraction from subsequent application.
@@
    Componential Analysis (CA): Parallels analytical decomposition Gluck et al. (2008), requiring models to reason over structured textual descriptions of individual images.
By comparing VLM performance across these paradigms, we can gain insights into their preferred processing modes and identify specific bottlenecks.
Critically, our framework, particularly the CA paradigm, offers unique advantages for VLM evaluation. It provides a mechanism to assess performance on inherently multi-image tasks (like BPs) even for VLMs primarily designed for single-image inputs. By transforming visual inputs into textual descriptions, CA allows us to disentangle the contributions of visual perception from higher-level reasoning. Providing high-quality, externally generated descriptions enables us to probe reasoning abilities largely independent of the model’s own perceptual fidelity, thereby isolating failures in visual processing versus failures in logical inference. This approach even permits evaluating the visual reasoning potential of uni-modal (text-only) LLMs.
Applying this framework to a range of contemporary VLMs, we find that leading models (GPT-4o, Gemini 2.0) exhibit strong performance, particularly when reasoning is structured (DRL, CA) or when based on high-quality textual descriptions (CA). However, a significant finding is the pronounced struggle of many open-source VLMs with the initial visual perception stage, as revealed by their poor performance in CA when using self-generated descriptions, which contrasts sharply with improved performance when perception is supported (see Ablations, Section 7). This suggests a critical perception bottleneck hindering downstream reasoning.
Our contributions are thus: (1) A novel, cognitively-inspired framework for the diagnostic evaluation of VLM visual reasoning. (2) Three evaluation paradigms enabling analysis of different reasoning strategies and facilitating multi-image task evaluation for diverse architectures. (3) A method (CA) for helping to disentangle perception from reasoning in VLMs and extending evaluation to LLMs. (4) Empirical results highlighting both the potential of advanced VLMs and a critical perception bottleneck in many current open-source models. (5) Evidence supporting the benefit of structured, multi-stage processing for VLM visual reasoning.
@@ 2 Related Work
Evaluating the burgeoning capabilities of VLMs necessitates diverse and challenging benchmarks. While foundational tasks like captioning Lin et al. (2014), VQA Antol et al. (2015); Gurari et al. (2018), and OCR Mishra et al. (2019) remain important, the focus has shifted towards more complex reasoning Zhang et al. (2024); Lu et al. (2022) and multi-modal understanding Meng et al. (2024). Benchmarks leveraging interleaved image-text data Zhu et al. (2024); Laurençon et al. (2024); Lin et al. (2024) or specifically designed for multi-image contexts Wu et al. (2023); Wang et al. (2024); Li et al. (2023) probe richer interactions. However, many benchmarks rely heavily on linguistic cues or do not isolate the type of abstract visual concept formation central to tasks like Bongard Problems.
Purely visual reasoning benchmarks force models to rely solely on visual input, providing a clearer window into their perceptual and non-verbal reasoning abilities. Bongard Problems Bongard (1968) are a classic example, requiring few-shot learning of abstract visual rules. Natural image variants like Bongard-HOI Jiang et al. (2022) and Bongard-OpenWorld Wu et al. (2024) increase ecological validity. We utilize these BPs as they demand comparison across multiple complex images to derive an underlying principle – a process akin to human concept learning. Other significant visual reasoning benchmarks targeting spatial, compositional, or abstract abilities include RPMs Barrett et al. (2018); Zhang et al. (2019), SVRT Fleuret et al. (2011), CVR Zerroug et al. (2022), V-PROM Teney et al. (2020), and ARC Chollet (2019). Our work contributes a methodology for analyzing how models solve such tasks.
Our evaluation framework is explicitly grounded in cognitive science. Human problem-solving employs diverse strategies Newell et al. (1972). Our paradigms are designed to reflect this diversity:
  • DVRL evokes holistic, gestalt-like processing or similarity-based judgments common in initial visual analysis Biederman (1987); von Helversen et al. (2014).
  • DRL mirrors deliberative, rule-based deductive reasoning where principles are explicitly formulated and applied Rips (1994); Knauff and Wolf (2010).
  • CA reflects analytical decomposition, where complex stimuli are broken into components for systematic analysis Gluck et al. (2008); Hutchinson (1992), related to feature-based learning.
By assessing VLMs through these lenses, we can compare their operational characteristics to known human cognitive modes, offering insights into the nature of their internal representations and computations.
Our work also resonates with research on eliciting structured reasoning in language models, notably Chain-of-Thought (CoT) prompting Wei et al. (2022); Kojima et al. (2022). CoT encourages models to articulate intermediate steps, often improving performance on complex tasks. While multimodal CoT is an active area Zhang et al. (2023); Chen et al. (2024), our paradigms differ. They represent distinct, cognitively-motivated evaluation conditions that structure the task itself (e.g., separating rule extraction, forcing textual representation) rather than solely prompting for verbalized steps within a single inference pass. We use these paradigms diagnostically to understand the impact of such structuring on visual reasoning (see Appendix A.6.6 for a comparison with CoT prompting). Our focus remains firmly on VLM evaluation and understanding the perception-reasoning interplay, distinct from areas like active learning.
@@ 3 Models
We selected a diverse suite of contemporary VLMs for evaluation, encompassing both leading closed-source systems and widely used open-source models of varying scales. This includes: GPT-4o OpenAI (2024), Gemini 2.0 Google (2024), Pixtral-12B Agrawal et al. (2024), Llama-Vision-3.2 (11B, 90B) Meta (2024), LLaVA (Llama-2 based; 7B, 13B, 34B) Liu et al. (2023), and LLaVA-Llama3-8B XTuner (2025). For ablation studies involving reasoning over text (Section 7.2), we also included text-only LLMs (e.g., Phi-4 Abdin et al. (2024); Qwen2.5 Yang et al. (2024); Deepseek-r1 Guo et al. (2025); Gemma2 Team et al. (2024)). Model access was via official APIs or the Ollama framework Ollama (2025) running locally on NVIDIA GPUs (2080Ti, 3090, 6000 Ada). All evaluations employed few-shot prompting with zero temperature for deterministic output; no model fine-tuning was performed. Appendix A.4 provides comprehensive details on models, configurations, etc.
@@ 4 Dataset and Task
Our primary testbed is the Bongard-OpenWorld dataset Wu et al. (2024). We selected this benchmark for its use of natural, real-world images and its requirement for few-shot abstract reasoning based on commonsense visual concepts, aligning with the cognitive challenges inherent in BPs. From the full dataset, we constructed a balanced subset of 500 test cases (see Appendix A.3 for sampling details). Each case comprises 6 positive images exemplifying a rule, 6 negative images violating it, and 1 query image for classification. The dataset’s grounding in real-world imagery and abstract rules provides a demanding test for VLM visual intelligence. An example task is shown in Figure 1. For generalizability assessment, we also utilize the Bongard-HOI Jiang et al. (2022) and Bongard Logo Nie et al. (2020) datasets (Section 6.2).
@@ 6.1 Performance on Bongard-OpenWorld
Componential Analysis (CA), reasoning over textual descriptions, yielded the highest accuracies for the top models (GPT-4o: 92.8%, Gemini 2.0: 93.6%), surpassing the reported human average Wu et al. (2024) for this benchmark. This underscores their powerful text-based reasoning ability when the visual input is effectively encoded into language. Pixtral-12B also achieved strong CA performance (87.2%). However, a significant gap emerged with other open-source models. Models like Llama-Vision and LLaVA variants exhibited much lower CA accuracy, often with dramatic imbalances between positive and negative sample performance (e.g., Llava-Llama3-8B: 53.2% overall, heavily biased towards negative samples). This pattern strongly suggests that the bottleneck for these models is not necessarily the abstract reasoning itself, but rather the fidelity of their internal visual perception and subsequent translation into usable representations (here, text descriptions).
@@
6.2 Generalizability Across Datasets
To probe the robustness of our findings and framework, we conducted evaluations on two additional datasets: Bongard-HOI Jiang et al. (2022) (natural images, human-object interactions) and the synthetic Bongard Logo Nie et al. (2020).
On Bongard-HOI, we evaluated GPT-4o and Gemini 2.0 across the four standard test splits (sosa, soua, uosa, uoua; N=100 each). The results, shown in Table 1, largely replicated the trends observed on Bongard-OpenWorld. Performance systematically improved with increased paradigm structure (DVRL < DRL < CA) for both models. CA yielded the highest accuracies (65-83% range), generally surpassing DRL (64-77%) and DVRL (61-75%). This consistency validates the framework’s applicability and the benefit of structured evaluation across different complex natural image reasoning datasets. Notably, model performance is lower overall on HOI compared to OpenWorld, suggesting HOI presents distinct challenges, potentially in identifying subtle interaction-based rules. The human average scores reported in Jiang et al. (2022) are high (87-95%), indicating a significant gap remains even for top models on this dataset.
@@ 6.2 Generalizability Across Datasets
Table 1: Performance (%) on Bongard-HOI splits across paradigms. Human average taken from Jiang et al. (2022) Splits: sosa: seen_obj_seen_act, soua: seen_obj_unseen_act, uosa: unseen_obj_seen_act, uoua: unseen_obj_unseen_act. Human average from cited source.
Evaluation on the synthetic Bongard Logo dataset yielded starkly different results. Both GPT-4o and Gemini 2.0 performed near chance level, achieving only around 50% accuracy. This dramatic drop compared to natural image datasets suggests significant difficulties for current VLMs in reasoning over abstract synthetic shapes. This could stem from biases in pre-training data favoring natural images or challenges inherent in the CA paradigm when trying to generate meaningful textual descriptions for abstract geometric forms.
These cross-dataset results indicate our framework is a useful tool for revealing performance patterns, showing consistent benefits of structure on natural images but also highlighting the distinct challenge posed by abstract synthetic reasoning tasks for contemporary VLMs.
Model DVRL (%) DRL (%) CA (%)
neg pos Overall neg pos Overall neg pos Overall
GPT-4o 66.4 93.6 80.0 82.8 93.2 88.0 92.8 92.8 92.8
Gemini 2.0 80.5 83.9 82.2 85.7 87.9 86.8 90.8 96.4 93.6
Pixtral-12B - - - - - - 88.8 85.6 87.2
Llama-Vision-11B - - - - - - 42.8 64.0 53.4
Llama-Vision-90B - - - - - - 58.7 51.5 55.1
Llava-7B - - - - - - 65.2 67.2 66.2
Llava-Llama3-8B - - - - - - 93.0 13.1 53.2
Human Average Wu et al. (2024) - - - - - - (across samples) 91.0

### v4

@@ Abstract
A fundamental challenge in artificial intelligence involves understanding the cognitive mechanisms underlying visual reasoning in sophisticated models like Vision-Language Models (VLMs). How do these models integrate visual perception with abstract thought, especially when reasoning across multiple images or requiring fine-grained compositional understanding? Drawing inspiration from cognitive science, this paper introduces a structured evaluation framework using diverse visual reasoning tasks – Bongard Problems (BPs) and Winoground – to dissect the perception-reasoning interface in VLMs. We propose three distinct evaluation paradigms, mirroring human problem-solving strategies: Direct Visual Rule Learning (DVRL; holistic processing), Deductive Rule Learning (DRL; rule extraction and application), and Componential Analysis (CA; analytical decomposition via task-agnostic textual descriptions). These paradigms systematically vary cognitive load and probe processing stages. Notably, CA enables multi-image reasoning evaluation even for single-image architectures and isolates reasoning from perception by operating on textual descriptions. Applying this framework, we demonstrate that CA, leveraging powerful language models for reasoning over rich, independently generated descriptions, achieves new state-of-the-art (SOTA) performance on challenging benchmarks including Bongard-OpenWorld, Bongard-HOI, and Winoground. Ablation studies confirm reasoning improves significantly when perceptual challenges are mitigated, revealing a critical perception bottleneck. Our framework provides a valuable diagnostic tool and suggests that decoupling perception (via rich, task-agnostic description) from reasoning is a promising direction for robust and general visual intelligence.
@@ 1 Introduction
Human cognition adeptly integrates visual perception with abstract reasoning to navigate and understand the world [1, 2]. A central goal in AI is to imbue machines with similar visual intelligence [3]. While Vision-Language Models (VLMs) show remarkable progress [4, 5], the underlying mechanisms enabling complex visual reasoning – especially tasks demanding abstraction across multiple images or fine-grained compositional understanding – remain opaque. How do VLMs bridge pixels and concepts, and how do their strategies compare to human cognition?
To investigate these questions, we utilize two distinct classes of challenging visual reasoning tasks. First, Bongard Problems (BPs) [6], a classic test requiring few-shot discovery of an abstract visual rule distinguishing positive and negative image sets. BPs demand perception, abstraction, and multi-image comparison, echoing human concept learning [7]. We use natural image variants, Bongard-OpenWorld [8] and Bongard-HOI [9]. Second, we use Winoground [10], which tests visio-linguistic compositional reasoning by requiring fine-grained differentiation between minimally contrastive image-caption pairs.
This paper introduces an evaluation framework explicitly designed to probe the cognitive processes underlying VLM performance on such challenging visual reasoning tasks, using natural image BPs [8] as our primary testbed. Our core contribution is an evaluation methodology grounded in cognitive science principles. We propose three distinct paradigms that structure the BP task differently, mirroring established human problem-solving strategies:
@@
    Direct Visual Rule Learning (DVRL): Simulates holistic or gist-based processing [11], requiring the model to analyze all images simultaneously.
@@
    Deductive Rule Learning (DRL): Mimics explicit, rule-based deduction [12], separating rule extraction from subsequent application.
@@
    Componential Analysis (CA): Parallels analytical decomposition [13], requiring models to reason over structured textual descriptions of individual images.
This framework allows systematic analysis of VLM behavior under different cognitive demands and identify specific bottlenecks.
The CA paradigm offers unique advantages. It facilitates multi-image reasoning evaluation even for single-image VLM architectures. Crucially, by generating comprehensive, task-agnostic image descriptions first and then performing reasoning solely over this text, CA allows us to disentangle perception from reasoning. Furthermore, by substituting externally generated high-quality descriptions, we can isolate reasoning capabilities from the model’s own perceptual limits and even evaluate text-only LLMs (Section 7.2).
Applying this framework, we achieve new state-of-the-art (SOTA) results on Bongard-OpenWorld, Bongard-HOI (on several splits), and Winoground using the CA paradigm, primarily by combining high-fidelity descriptions with powerful reasoning models (including GPT-4o, Gemini 2.0, and specialized LLMs). This success across diverse tasks suggests the robustness and generality of decoupling perception into rich textual representations and leveraging strong language-based reasoning. Concurrently, our analysis reveals a significant perception bottleneck in many open-source VLMs, whose performance drastically improves when their perceptual front-end is bypassed.
Our contributions are thus: (1) A novel, cognitively-inspired framework for the diagnostic evaluation of VLM visual reasoning. (2) Three evaluation paradigms enabling analysis of different reasoning strategies and facilitating multi-image task evaluation for diverse architectures. (3) A method (CA) for helping to disentangle perception from reasoning in VLMs and extending evaluation to LLMs. (4) Empirical results highlighting both the potential of advanced VLMs and a critical perception bottleneck in many current open-source models. (5) Evidence supporting the benefit of structured, multi-stage processing for VLM visual reasoning – SOTA performance on multiple benchmarks (Bongard-OW, HOI, Winoground)
@@ 2 Related Work
Evaluating the burgeoning capabilities of Vision-Language Models (VLMs) necessitates diverse benchmarks. While foundational tasks like Visual Question Answering (VQA) [14, 15] remain important, the focus has shifted towards complex reasoning [16, 17] and multi-modal, multi-image understanding, utilizing interleaved corpora [18, 19] and dedicated benchmarks [20, 21, 22]. However, many evaluations rely heavily on linguistic context or do not specifically isolate the abstract visual concept formation and relational reasoning central to purely visual challenges like Bongard Problems (BPs).
Purely visual reasoning benchmarks provide a clearer window into perceptual and non-verbal reasoning abilities. BPs [6] are a canonical example, testing few-shot abstract rule discovery. Natural image variants like Bongard-HOI [9] and Bongard-OpenWorld [8] enhance ecological validity, demanding comparison across complex images akin to human concept learning. Alongside other abstract reasoning tests like Raven’s Progressive Matrices (RPMs) [23, 24], BPs provide a crucial testbed for advanced visual intelligence. Winoground [10] specifically targets visio-linguistic compositional understanding. Our contribution here an evaluation methodology designed to analyze how models approach such cognitively demanding tasks.
Our evaluation framework is explicitly grounded in cognitive science perspectives on human problem-solving [25]. The proposed paradigms reflect diverse cognitive strategies: Direct Visual Rule Learning (DVRL) mirrors rapid holistic or similarity-based processing common in initial visual analysis [11, 26]; Deductive Rule Learning (DRL) reflects explicit, deliberative rule-based deduction [12, 27]; and Componential Analysis (CA) parallels analytical decomposition where problems are broken into constituent features or components for systematic reasoning [13, 28]. Assessing VLMs through these distinct lenses provides a structured way to probe their internal processing characteristics relative to known cognitive modes.
Our approach connects to Chain-of-Thought (CoT) prompting methods [29, 30] that structure reasoning, including recent multimodal extensions [31, 32] such as DDCoT [33], CoCoT [34], and Compositional CoT [35]. While sharing the goal of step-wise processing, a key distinction lies in our Componential Analysis (CA) paradigm: whereas many multimodal CoT approaches generate intermediate representations often conditioned on the specific task context (e.g., descriptions tailored to a given caption [33, 34]), CA deliberately generates comprehensive, task-independent textual image descriptions in a separate first stage, with subsequent reasoning operating only on this output. This decoupling appears key to the generalizability and SOTA performance we observe. Our paradigms act as diagnostic evaluation conditions manipulating task structure, rather than solely performance-boosting prompts. Consequently, our paradigms function primarily as distinct, cognitively-motivated evaluation conditions that manipulate the task structure itself (e.g., comparing holistic vs. deductive processing, separating perception from reasoning via CA) for diagnostic insights into VLM processes, rather than being prompting techniques aimed solely at maximizing task performance.
@@ 3 Models
We selected a diverse suite of contemporary VLMs for evaluation, encompassing both leading closed-source systems and widely used open-source models of varying scales. This includes: GPT-4o [36], Gemini 2.0 [37], Pixtral-12B [38], Llama-Vision-3.2 (11B, 90B) [39], LLaVA (Llama-2 based; 7B, 13B, 34B) [40], and LLaVA-Llama3-8B [41]. For ablation studies involving reasoning over text (Section 7.2), we also included text-only LLMs (e.g., Phi-4 [42]; Qwen2.5 [43]; Deepseek-r1 [44]; Gemma2 [45]). Model access was via official APIs or the Ollama framework [46] running locally on NVIDIA GPUs (2080Ti, 3090, 6000 Ada). All evaluations employed few-shot prompting with zero temperature for deterministic output; no model fine-tuning was performed. Appendix A.4 provides comprehensive details on models, configurations, etc.
@@ 4 Dataset and Task
Our primary testbed is the Bongard-OpenWorld dataset [8]. We selected this benchmark for its use of natural, real-world images and its requirement for few-shot abstract reasoning based on commonsense visual concepts, aligning with the cognitive challenges inherent in BPs. From the full dataset, we constructed a balanced subset of 500 test cases (see Appendix A.3 for sampling details). Each case comprises 6 positive images exemplifying a rule, 6 negative images violating it, and 1 query image for classification. The dataset’s grounding in real-world imagery and abstract rules provides a demanding test for VLM visual intelligence. An example task is shown in Figure 1. For generalizability assessment, we also utilize the Bongard-HOI [9] and Bongard Logo [47] datasets (Section 6.2).
We use additional two distinct visual reasoning benchmarks:
  • Bongard-HOI (BP-HOI) [9]: Focuses on rules based on human-object interactions in natural images. Used 4 standard splits (N=100 each).
  • Winoground [10]: Assesses fine-grained compositional visio-linguistic reasoning via image-caption matching (N=400 samples used).
These diverse tasks probe abstraction (BPs) and compositionality (Winoground), providing a comprehensive testbed.
@@ 6.1 Performance on Bongard-OpenWorld
Model DVRL (%) DRL (%) CA (%)
neg pos Overall neg pos Overall neg pos Overall
GPT-4o [36] 66.4 93.6 80.0 82.8 93.2 88.0 92.8 92.8 92.8
Gemini 2.0 [37] 80.5 83.9 82.2 85.7 87.9 86.8 90.8 96.4 93.6
Pixtral-12B - - - - - - 88.8 85.6 87.2
Llama-Vision-11B [39] - - - - - - 42.8 64.0 53.4
Llama-Vision-90B [39] - - - - - - 58.7 51.5 55.1
Llava-7B [40] - - - - - - 65.2 67.2 66.2
Llava-Llama3-8B [41] - - - - - - 93.0 13.1 53.2
Human Average [8] - - - - - - (across samples) 91.0
@@
Componential Analysis (CA), reasoning over textual descriptions, yielded the highest accuracies for the top models (GPT-4o: 92.8%, Gemini 2.0: 93.6%), surpassing the reported human average [8] for this benchmark. This underscores their powerful text-based reasoning ability when the visual input is effectively encoded into language. Pixtral-12B also achieved strong CA performance (87.2%). However, a significant gap emerged with other open-source models. Models like Llama-Vision and LLaVA variants exhibited much lower CA accuracy, often with dramatic imbalances between positive and negative sample performance (e.g., Llava-Llama3-8B: 53.2% overall, heavily biased towards negative samples). This pattern strongly suggests that the bottleneck for these models is not necessarily the abstract reasoning itself, but rather the fidelity of their internal visual perception and subsequent translation into usable representations (here, text descriptions).
@@
6.2 Performance on Bongard HOI
@@
Table 2: Performance (%) on Bongard-HOI splits across paradigms. Human average taken from [9] Splits: sosa: seen_obj_seen_act, soua: seen_obj_unseen_act, uosa: unseen_obj_seen_act, uoua: unseen_obj_unseen_act. Human average from cited source.
On Bongard-HOI, we evaluated GPT-4o and Gemini 2.0 across the four standard test splits (sosa, soua, uosa, uoua; N=100 each). The results, shown in Table 2, largely replicated the trends observed on Bongard-OpenWorld. Performance systematically improved with increased paradigm structure (DVRL < DRL < CA) for both models. CA yielded the highest accuracies (65-83% range), generally surpassing DRL (64-77%) and DVRL (61-75%). This consistency validates the framework’s applicability and the benefit of structured evaluation across different complex natural image reasoning datasets. Notably, model performance is lower overall on HOI compared to OpenWorld, suggesting HOI presents distinct challenges, potentially in identifying subtle interaction-based rules. The human average scores reported in [9] are high (87-95%), indicating a significant gap remains even for top models on this dataset.
6.3 Performance on Winoground
Model Text Score Image Score Group Score
GPT-4o [36] 75.50 58.50 52.00
Gemini 2.0 [37] 71.00 48.75 42.00
Llama3.3-70B ([48]) 68.25 49.25 41.75
Qwen2.5-32B ([43]) 67.00 46.25 40.00
Phi-4-14B ([42]) 65.25 46.00 37.75
Qwen2.5-14B ([43]) 59.25 34.50 27.25
MMICL + CoCoT [34] 64.25 52.5 50.75
Table 3: State-of-the-art performance on the Winoground benchmark achieved using our Componential Analysis (CA) paradigm. Scores reported are the standard Winoground metrics: Text Score (correct caption selection per image description), Image Score (correct image description selection per caption), and Group Score (all selections correct per sample), averaged over 400 samples.
We applied our CA paradigm to the distinct challenge of Winoground [10], which requires fine-grained compositional reasoning to match images and captions. Stage 1 generated task-agnostic descriptions for each image using Gemini 2.0, and Stage 2 used various LLMs (acting as the reasoning engine) to perform the matching based only on the descriptions and captions (see Appendix A.4.4 for metric details).
As shown in Table 3, this approach achieves new state-of-the-art results on Winoground across all three metrics (Text, Image, Group scores). Using GPT-4o as the reasoning engine in CA yields scores of Text: 75.5%, Image: 58.5%, Group: 52.0%. This significantly surpasses previously reported SOTA. Even using Gemini 2.0 or strong open LLMs like Llama3-70B or Qwen2.5-32B within the CA framework produces highly competitive or SOTA results.
This successful application to Winoground demonstrates that our CA evaluation strategy is not limited to rule-discovery tasks like Bongard problems but can be effectively adapted to probe compositional understanding, further supporting the generalizability and diagnostic potential of our cognitively-inspired framework. This success is significant because our CA approach generates image descriptions independently of the captions or the matching task itself. Unlike prior methods that often condition visual analysis on textual context [33, 34], our task-agnostic description generation followed by powerful text-based reasoning proves highly effective for fine-grained compositionality. This suggests that decoupling rich perception (into text) from robust linguistic reasoning offers a generalizable and high-performing strategy for complex visio-linguistic tasks.